Taronne received his cell phone bill in the mail. The bill was for $99.00 and the sales tax was 4.5%. How much was he charged in taxes?
1
Expert's answer
2020-10-27T18:59:28-0400

Let the origin bill without tax is $x.

He was charged $ x*"\dfrac{4.5}{100}" in tax.


Therefore x+x*"\dfrac{4.5}{100}=99"


"\Rightarrow x=94.73"


He was charged in taxes=(94.73 *"\dfrac{4.5}{100})= 4.26" in dollars